OpenStudy (anonymous):

PLEASE HELP! I need the work shown please. x + 2 + x = 22

OpenStudy (anonymous):

x + x + 2 = 22 2x + 2 = 22 2x = 20 x = 20/2 = 10

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (anonymous):

ur welcome